AZ Products has 140,000 shares of common stock outstanding at a market price of $27 a

share. AZ Products just paid an annual dividend of $1.25 a share and the dividend growth rate is

2 percent. The company also has 2,500 bonds outstanding with a face value of $1,000 per bond.

The bonds have a pretax yield of 7.35 percent and sell at 108.2 percent of face value. The

company's tax rate is 21 percent. What is the weighted average cost of capital?

=6.34%

-Please check the answers and show all work typed out. No excel or grid style please as I am on mobile.
Answer #1

We know that WACC = Weight of equity * cost of equity + weight of debt * post tax cost of debt.

Equity

Cost of equity = (Expected dividend next year / Price of share today ) + Growth rate

Expected dividend next year = 1.25*102% = 1.275

therefore cost of equity = (1.275/27)+2% = 6.72%

Value of equity = 140000*27 = 3780000

Debt

Post tax cost of debt = YTM*(1-tax rate) = 7.35%*(1-21%) = 5.81%

Value of debt = 1000*108.2%*2500 = 2705000

Computation of WACC

WACC = (5.81%*2705000+6.72%*3780000)/(2705000+3780000)= 6.34%